A camera with a 89-mm focal length lens is being used to photograph a person standing 4.68 m away. How far from the lens must the film be (in m)?

answer to 3 decimal places


1
Expert's answer
2021-06-07T08:31:49-0400

"\frac 1F=\frac 1d+\frac 1f\implies" "d=\frac{Ff}{f-F},"

"d=\frac{0.089\cdot 4.68}{4.68-0.089}=0.091~m."